2. Tharaldson Corporation makes a product with the following standard costs: Unit Standard Quantity or Hours 5.7 ounce...

Tharaldson Corporation makes a product with the following standard costs: 


Standard
Standard

Quantity orStandard Price orCost Per

HoursRateUnit
Direct materials5.7 ounces$2.00 per ounce$11.40
Direct labor0.8 hours$11.00 per hour$8.80
Variable overhead0.8 hours$6.00 per hour$4.80

The company reported the following results concerning this product in June.

Originally budgeted output3,900 units
Actual output3,500 units
Raw materials used in production20,700 ounces
Purchases of raw materials21,800 ounces
Actual direct labor-hours6,000 hours
Actual cost of raw materials purchases$42,600
Actual direct labor cost$13,900
Actual variable overhead cost$3,950

The company applies variable overhead on the basis of direct labor-hours. The direct materials purchases variance is computed when the materials are purchased. 

The materials price variance for June is:

2.

Milar Corporation makes a product with the following standard costs: 


Standard Quantity or HoursStandard Price or Rate



Direct materials9.5 pounds$9.00 per pound
Direct labor0.8 hours$31.00 per hour
Variable overhead0.8 hours$14.50 per hour


The company applies variable overhead on the basis of direct labor-hours. The direct materials purchases variance is computed when the materials are purchased. In January the company produced 3,420 units using 13,680 pounds of the direct material and 2,856 direct labor-hours. During the month, the company purchased 14,440 pounds of the direct material at a cost of $35,100. The actual direct labor cost was $88,125 and the actual variable overhead cost was $39,700. 

The labor rate variance for January is:


Kartman Corporation makes a product with the following standard costs: 


Standard Quantity or HoursStandard Price or RateStandard Cost Per Unit




Direct materials7.6 pounds$8.10 per pound$61.56
Direct labor0.6 hours$35.00 per hour$21.00
Variable overhead0.6 hours$5.10 per hour$3.06


In June the company's budgeted production was 4,500 units but the actual production was 4,600 units. The company used 23,250 pounds of the direct material and 2,400 direct labor-hours to produce this output. During the month, the company purchased 26,500 pounds of the direct material at a cost of $181,180. The actual direct labor cost was $58,121 and the actual variable overhead cost was $11,661. 

The company applies variable overhead on the basis of direct labor-hours. The direct materials purchases variance is computed when the materials are purchased. 

The variable overhead rate variance for June is:
Answer #1

1. Materials Price Variance

= (SP-AP) * AQ

= (2 - 42,600/21,800) * 21,800

= (2-1.95) * 21,800

= 1000 Favourable